# A Survey of Deep Graph Learning under Distribution Shifts: from Graph Out-of-Distribution Generalization to Adaptation

### v1

@@ Abstract
Distribution shifts on graphs – the discrepancies in data distribution between training and employing a graph machine learning model – are ubiquitous and often unavoidable in real-world scenarios. These shifts may severely deteriorate model performance, posing significant challenges for reliable graph machine learning. Consequently, there has been a surge in research on graph machine learning under distribution shifts, aiming to train models to achieve satisfactory performance on out-of-distribution (OOD) test data. In our survey, we provide an up-to-date and forward-looking review of deep graph learning under distribution shifts. Specifically, we cover three primary scenarios: graph OOD generalization, training-time graph OOD adaptation, and test-time graph OOD adaptation. We begin by formally formulating the problems and discussing various types of distribution shifts that can affect graph learning, such as covariate shifts and concept shifts. To provide a better understanding of the literature, we systematically categorize the existing models based on our proposed taxonomy and investigate the adopted techniques behind. We also summarize commonly used datasets in this research area to facilitate further investigation. Finally, we point out promising research directions and the corresponding challenges to encourage further study in this vital domain. Additionally, we provide a continuously updated reading list at https://github.com/kaize0409/Awesome-Graph-OOD.
@@ 1 Introduction
While graph machine learning has achieved notable success, most of the existing efforts assume that test data follows the same distribution as training data, which is often invalid in the wild. When confronted with Out-Of-Distribution (OOD) samples, the performance of graph machine learning methods may substantially degrade, limiting their efficacy in high-stake graph applications such as finance and healthcare [7]. Although numerous transfer learning methods have been proposed to address distribution shifts for Euclidean data [8, 9, 10], their direct application to graph data is challenging. This is due to the interconnected nature of entities on graphs, which violates the independent and identically distributed (IID) assumption inherent in traditional transfer learning methods. Moreover, the various types of graph shifts introduce new challenges. These shifts occur across different modalities including features, structures, and labels, and can manifest in various forms such as variations in graph sizes, subgraph densities, and homophily [11]. Given these obstacles, increasing research efforts have been dedicated to improving the reliability of graph machine learning against distribution shifts, concentrating on three main scenarios: graph OOD generalization  [7, 11], training-time graph OOD adaptation [12, 13], and test-time graph OOD adaptation [14, 15].
@@
The primary distinction between graph OOD generalization and adaptation methods lies in their assumptions regarding the availability of target data. Graph OOD generalization methods typically assume the unavailability of target data during model training and aim to enhance the model’s generalization performance on any potential unseen test distribution. In contrast, both training-time and test-time adaptation methods assume the availability of target data and aim to improve model performance on this specific target. However, they differ in their assumptions about the source data and in how they utilize knowledge of the source distribution. Training-time adaptation assumes that both the source and target graphs are available simultaneously, allowing model adaptation to start from scratch during the training process. On the other hand, test-time adaptation typically assumes access to a model pre-trained on the source graph, rather than the source graph itself, and begins adapting the model to the target data from this pre-trained state. Although graph OOD generalization, training-time OOD adaptation, and test-time OOD adaptation are closely related, there is currently no unified framework that comprehensively discusses deep graph learning under distribution shifts across all three scenarios.
@@
With recent progress on graph OOD learning, an up-to-date and forward-looking review of this field is urgently needed. In this survey, we provide, to the best of our knowledge, the first unified and systematic review of the literature on deep graph learning under distribution shifts. We start by formally formulating the problems and discussing different types of graph distribution shifts in graph machine learning. Next, our new taxonomy is proposed, classifying existing methods into three categories based on the model learning scenario: (1) graph OOD generalization, where the generalizability is enhanced through strategic design of the model when training on source data, (2) training-time graph OOD adaptation, where the adaptation is performed when jointly training the model based on both source and target data [16, 17], and (3) test-time graph OOD adaptation, where the adaptation happens when adjusting a pre-trained source model to the target data [18, 19].
@@
To deepen our understanding of these approaches, we further classify existing methods within each of the three categories into model-centric and data-centric strategies. Model-centric approaches focus on the learning process or the architecture of the graph model itself, enhancing the model’s inherent ability to generalize or adapt to distribution shifts by refining its structure, training objectives, or learning mechanisms. In contrast, data-centric approaches emphasize the manipulation of input graphs, improving model performance by addressing the data directly, either through preprocessing techniques or data augmentation strategies. Within each subline of research, we elaborate on the detailed techniques for enhancing the generalizability or adaptability under distribution shifts on graphs. Additionally, we provide a summary of the datasets utilized in these studies, highlighting their characteristics and relevance to the challenges posed by distribution shifts. Based on the current progress on graph OOD learning, at the end we also point out several promising research directions in this evolving field.
@@
Differences between this survey and existing ones. Despite the urgent need for an overview of graph learning under distribution shifts, existing surveys have primarily focused on subfields within this area, rather than providing a comprehensive overview from multiple scenarios. Until now, there have been several surveys in related areas, but exhibiting distinct focuses, including graph OOD generalization [7, 20], graph domain adaptation [21, 22], trustworthy graph learning [23] related to distribution shifts. Our work distinguishes itself from existing ones in the following aspects: (1) Main focus. Our survey centers on the challenges and solutions for graph learning under distribution shifts, while [23] analyzes OOD issues from a trustworthy perspective and does not delve into the methodological aspects. Conversely, [20] examines graph machine learning from a causal perspective, which is narrower than our broad examination. (2) Taxonomy. We provide a comprehensive categorization of existing methods and summarize them, whereas related work, such as [24], lacks such summaries. Other surveys like [21] and [22] primarily focus on domain adaptation without addressing the broader scope of graph OOD learning. Additionally, we provide coverage of the most recent advancements and discussions in this field.
@@
Survey structure. The general organization of this survey is presented as follows: Section 2 introduces the notations and preliminaries. Sections 3, 4 and 5 review graph OOD generalization, training-time graph OOD adaptation, and test-time graph OOD adaptation, respectively. Each section discusses model-centric and data-centric approaches within its scenario, further detailing the techniques associated with each category. Furthermore, Section 6 provides a comprehensive summary of the datasets used in the literature, highlighting popular graph datasets for evaluation and their relevance to the challenges posed by distribution shifts. Section 7 explores promising future research directions and the associated challenges. Finally, Section 8 presents the conclusion of this survey.
2 Preliminaries
2.1 Problem Definition
Figure 1: An illustration of graph OOD generalization, training-time graph OOD adaptation, and test-time graph OOD adaptation. The source graph can be single or multiple.
@@ 2.1 Problem Definition
In this section, we take node classification as an exemplar to define node-level tasks of graph learning under distribution shifts. We consider three distinct problem scenarios (i.e., Graph OOD Generalization, Training-time Graph OOD Adaptation, and Test-time Graph OOD Adaptation) and the illustration can be found in Figure 1. Note that these definitions can be readily extended to encompass edge-level and graph-level learning tasks. Specifically, let $\varphi_{\theta}$ denote a graph model characterized by parameters $\theta$. Generally, the deep graph model $\varphi$ can be decomposed as $g\circ f$, where $g(\cdot):(\textbf{A},\textbf{X})\rightarrow\textbf{H}$ is a graph encoder mapping the graph inputs to latent representations H, and $f(\cdot)$ is a classifier in the latent space.
@@
Graph OOD Generalization: In this scenario, we assume that distribution shifts exist between source and target distributions $\mathcal{P}_{S}\neq\mathcal{P}_{T}$, and that $\mathcal{P}_{T}$ is not observable during model training. The objective of node-level graph OOD generalization is to learn an optimal model $\varphi_{{\theta}^{*}}$ using the nodes and labels from the source graph $(\mathcal{G}_{S},\textbf{Y}_{S})$, such that it generalizes well to the unseen target graph $\mathcal{G_{T}}$.
@@
If $\textbf{Y}_{T}=\varnothing$, indicating that none of the nodes in the target graph are labeled, the problem is classified as unsupervised domain adaptation. The learning objective $l(\cdot)$ exploits source domain labels but would not rely on target domain labels. Otherwise, when a portion of the nodes in the target graph are labeled, the problem is considered semi-supervised, and $l(\cdot)$ can utilize both the source domain labels and the partial target domain labels to optimize model performance.
@@
Test-time Graph OOD Adaptation: In this scenario, a pre-trained model $\varphi_{\theta^{\prime}}$ and a target graph $\mathcal{G_{T}}$ with node labels $\textbf{Y}_{T}$ (which may be partially labeled or entirely unlabeled) are given. The goal is to adapt the pre-trained model during the test-time training phase in response to distribution shifts in the target domain. This adaptation is performed without access to source domain data during this phase. Given the distribution shifts between the source and target domains, the objective is to adjust the pre-trained model $\varphi_{\theta^{\prime}}$ at test time to optimize its performance on the target graph:
@@
Here, $\theta^{\prime}$ represents the parameters from the pre-trained model, and $\theta$ represents the parameters introduced for test-time adaptation.
Graph Distribution Shift Types. In traditional machine learning, several studies have discussed and defined various types of distribution shifts [25, 26], of which the most widely-used concepts are covariate shifts (i.e., shifts in $\mathcal{P}(\textbf{X})$) and concept shifts (i.e., shifts in $\mathcal{P}(\textbf{X}|\textbf{Y})$ or $\mathcal{P}(\textbf{Y}|\textbf{X})$). These concepts can naturally be extended to graph setting by replacing feature inputs X with the graph inputs $\mathcal{G}=(\textbf{A},\textbf{X})$.
@@
2.2 Discussion on Related Topics
Several topics are closely related to graph OOD learning, including: graph transfer learning, graph domain adaptation, graph OOD detection, and graph continual learning. These topics, sharing similar goals with graph OOD learning but exhibiting nuanced differences, are discussed as follows.
Graph Transfer Learning. Traditionally, transfer learning is a more wide-ranging concept and involves transferring knowledge across distribution changes and diverse tasks. Compared to the focus of this survey – graph learning under distribution shifts, graph transfer learning does not exclusively focus on addressing distribution shifts, but also includes leveraging knowledge acquired from one task to enhance performance in another context.
Graph Domain Adaptation. Following traditional domain adaptation, graph domain adaptation methods [21, 22] typically rely on the covariate shift assumption about an invariant relationship between graph inputs and labels, represented as $\mathcal{P}_{S}(\textbf{Y}|\textbf{A},\textbf{X})=\mathcal{P}_{T}(\textbf{Y}|\textbf{A},\textbf{X})$. The objective is then to address distribution shifts in the input graph space across domains $\mathcal{P}_{S}(\textbf{A},\textbf{X})\neq\mathcal{P}_{T}(\textbf{A},\textbf{X})$. In contrast to graph domain adaptation, graph OOD learning is more general and comprehensive, involving distribution shifts that go beyond the covariate shift assumption.
@@ 2.2 Discussion on Related Topics
2.3 Proposed Taxonomy
As seen from the problem definitions, graph OOD generalization, training-time graph OOD adaptation, and test-time graph OOD adaptation, significantly differ in their model learning scenarios. In particular, graph OOD generalization methods assume the inaccessibility of target data and involve only source data; training-time graph OOD adaptation methods involve both source and target data and start model training from scratch, while test-time graph OOD adaptation methods assume the unavailability of source data and aim to adapt a pre-trained source model to the target. Accordingly, in the following three sections, we firstly categorize existing methods into graph OOD generalization, training-time graph OOD adaptation, and test-time graph OOD adaptation. Within each section, we further classify methods into model-centric and data-centric approaches. Model-centric approaches center on the learning process or the design of the graph model, while data-centric approaches emphasize the manipulation of input graphs, such as transforming graph structures or features. Our taxonomy is illstrated in Figure 2.
3 Graph OOD Generalization
In the field of graph machine learning against distribution shifts, the scenario of graph OOD generalization problems has attracted increasing attention in recent years [33, 34, 35, 36]. The setting of graph OOD generalization assumes that the model has no access to the target domain during training, which is notably more challenging than the domain adaptation setting, where the target domain is known in advance [5, 18]. Therefore, it is particularly crucial to enhance the overall generalizability of the model, such that it can handle samples from any unseen domain [37]. In this section, we provide a comprehensive review of the most recent advances in graph OOD generalization, focusing on the model-centric and data-centric approaches. A summary of existing methods is presented in Table I.
3.1 Model-Centric Approaches

### v2

@@ Abstract
Distribution shifts on graphs – the discrepancies in data distribution between training and employing a graph machine learning model – are ubiquitous and often unavoidable in real-world scenarios. These shifts may severely deteriorate model performance, posing significant challenges for reliable graph machine learning. Consequently, there has been a surge in research on graph machine learning under distribution shifts, aiming to train models to achieve satisfactory performance on out-of-distribution (OOD) test data. In our survey, we provide an up-to-date and forward-looking review of deep graph learning under distribution shifts. Specifically, we cover three primary scenarios: graph OOD generalization, training-time graph OOD adaptation, and test-time graph OOD adaptation. We begin by formally formulating the problems and discussing various types of distribution shifts that can affect graph learning, such as covariate shifts and concept shifts. To provide a better understanding of the literature, we introduce a systematic taxonomy that classifies existing methods into model-centric and data-centric approaches, investigating the techniques used in each category. We also summarize commonly used datasets in this research area to facilitate further investigation. Finally, we point out promising research directions and the corresponding challenges to encourage further study in this vital domain. We also provide a continuously updated reading list at https://github.com/kaize0409/Awesome-Graph-OOD.
@@ I Introduction
While graph machine learning has achieved notable success, most of the existing efforts assume that test data follows the same distribution as training data, which is often invalid in the wild. When confronted with Out-Of-Distribution (OOD) samples, the performance of graph machine learning methods may substantially degrade, limiting their efficacy in high-stake graph applications such as finance and healthcare [7]. Although numerous transfer learning methods have been proposed to address distribution shifts for Euclidean data [8, 9, 10], their direct application to graph data is challenging. This is due to the interconnected nature of entities on graphs, which violates the independent and identically distributed (IID) assumption inherent in traditional transfer learning methods. Moreover, the various types of graph shifts introduce new challenges. These shifts occur across different aspects including features, structures, and labels, and can manifest in various forms such as variations in graph sizes, subgraph densities, and homophily [11]. Given these obstacles, increasing research efforts have been dedicated to improving the reliability of graph machine learning against distribution shifts, concentrating on three main scenarios: graph OOD generalization  [7, 11], training-time graph OOD adaptation [12, 13], and test-time graph OOD adaptation [14, 15].
@@
The primary distinction between graph OOD generalization and adaptation methods lies in their assumptions regarding the availability of target data. Graph OOD generalization assumes the unavailability of target data during model training and aims to enhance the model’s generalization performance on potentially unseen test distributions. In contrast, both training-time and test-time adaptation assume the availability of target data and focus on improving model performance for a specific target domain. However, they differ in their use of source data and how they leverage source distribution knowledge. Training-time adaptation assumes that both the source and target graphs are available simultaneously, allowing the model adaptation to start from scratch during the training process. On the other hand, test-time adaptation typically assumes access to a model pre-trained on the source domain, rather than the source graph itself, and begins adapting the model to the target data from this pre-trained state. Although graph OOD generalization, training-time OOD adaptation, and test-time OOD adaptation are closely related, there is currently no unified framework that comprehensively discusses deep graph learning under distribution shifts across all three scenarios.
@@
With recent progress on graph OOD learning, an up-to-date and forward-looking review of this field is urgently needed. In this survey, we provide, to the best of our knowledge, the first unified and systematic review of the literature on deep graph learning under distribution shifts. We start by formally formulating the problems and discussing different types of graph distribution shifts in graph machine learning. Next, our new taxonomy is proposed, classifying existing methods into three categories based on the model learning scenario: (1) graph OOD generalization, which typically relies on strong regularization techniques and domain-invariant features to generalize well to unseen domains, (2) training-time graph OOD adaptation, which offers a more flexible approach by allowing access to both source and target data, making it more suitable for situations where domain shift is expected but some target data is available [16, 17], and (3) test-time graph OOD adaptation, which does not have access to the source data during adaptation and focuses on leveraging a pre-trained source model to adjust to the target domain at test time. This approach is especially useful in online learning or real-time adaptation scenarios [18, 19].
@@
To deepen our understanding of these approaches, we further classify existing methods within each of the three scenarios into model-centric and data-centric strategies. Model-centric approaches focus on improving the learning process or designing the architecture of the graph model. These methods enhance the model’s ability to generalize or adapt to distribution shifts by refining key aspects such as network structure, training objectives, or learning mechanisms. On the other hand, data-centric approaches emphasize the manipulation of input graphs, improving model performance by addressing the data directly, either through preprocessing techniques or data augmentation strategies. Within each subfield of research, we systematically analyze the detailed techniques designed to enhance either the generalizability or adaptability of graph models under distribution shifts. Furthermore, we provide a summary of the datasets utilized in these studies, highlighting their characteristics and relevance to the challenges posed by distribution shifts. Despite the progress in graph OOD generalization and adaptation, several open challenges remain. In the end we also point out several promising research directions in this evolving field.
@@
Differences between this survey and existing ones. There is a growing need for a comprehensive review of graph learning under distribution shifts, while existing surveys have primarily focused on specific subdomains rather than providing a unified perspective across different scenarios. Until now, several related surveys have been conducted, each with a distinct focus, including graph OOD generalization [7, 20], graph domain adaptation [21, 22], trustworthy graph learning [23] related to distribution shifts. Our work differs from these surveys in the following key aspects: (1) Main focus. We provide an in-depth analysis of both challenges and methodological advancements in graph learning under distribution shifts. In contrast, [23] analyzes OOD issues from a trustworthy perspective but does not delve into the methodological aspects. Similarly, [20] approaches the problem through a causal lens, which, while valuable, offers a narrower perspective compared to our broader investigation. (2) Taxonomy. We introduce a comprehensive categorization of existing methods and systematically summarize them. In comparison, prior surveys such as [24] lack a comprehensive taxonomy, while [21] and [22] primarily concentrate on domain adaptation, without addressing the broader landscape of graph OOD learning. Additionally, we provide coverage of the most recent advancements and discussions in this field, capturing emerging trends, newly proposed techniques, and ongoing challenges.
@@
Survey structure. The general organization of this survey is presented as follows: Section II introduces the notations and preliminaries. Sections III, IV and V review graph OOD generalization, training-time graph OOD adaptation, and test-time graph OOD adaptation, respectively. Each section discusses model-centric and data-centric approaches within its scenario, further detailing the techniques associated with each category. Section VI provides a comprehensive summary of the datasets used in the literature, highlighting popular graph datasets for evaluation and their relevance to the challenges posed by distribution shifts. Section VII explores promising future research directions and the associated challenges. Finally, Section VIII presents the conclusion of this survey.
II Preliminaries
II-A Problem Definition
@@
In this section, we take node classification as an exemplar to define node-level tasks of graph learning under distribution shifts. However, these definitions can be readily extended to other tasks, such as edge-level and graph-level learning tasks, which are also discussed throughout the survey. Specifically, let $\varphi_{\theta}$ denote a graph model characterized by parameters $\theta$. Generally, the deep graph model $\varphi$ can be decomposed as $g\circ f$, where $g(\cdot):(\textbf{A},\textbf{X})\rightarrow\textbf{H}$ is a graph encoder mapping the graph inputs to latent representations H, and $f(\cdot)$ is a classifier in the latent space. We summarize the notations in Table I.
TABLE I: Frequently used notations.
Notation Description
$\mathcal{G}=(\textbf{A},\textbf{X})$ The (original) graph.
$\mathcal{V}$ The set of nodes.
A The adjacency matrix.
X Node feature matrix.
Y Node label matrix.
H The latent representations of the graph.
$K$ The number of classes in the node classification task.
$N$ The number of nodes.
$\mathcal{P}_{S},\mathcal{P}_{T}$ Source and target distributions.
$\varphi_{\theta}$ The graph model characterized by parameters $\theta$.
$g(\cdot)$ The graph encoder function.
$g_{s}(\cdot)$ The graph encoder that extracts invariant task-related information.
$g_{o}(\cdot)$ The graph encoder used for other components, excluding $g_{s}$.
$f(\cdot)$ The classifier function.
$f_{d}(\cdot)$ The domain discriminator function.
$l(\cdot)$ The learning objective.
We consider three distinct problem scenarios (i.e., Graph OOD Generalization, Training-time Graph OOD Adaptation, and Test-time Graph OOD Adaptation) and the illustration can be found in Figure 1.
Figure 1: An illustration of the distinctions between graph OOD generalization, training-time graph OOD adaptation, and test-time graph OOD adaptation. The primary difference lies in the availability of target data and how source data is utilized. Graph OOD generalization assumes no target data during training, while both training-time and test-time adaptations assume access to target data. Training-time adaptation uses both source and target data during training, whereas test-time adaptation adapts a pre-trained model with only target data. The source graph can be single or multiple.
@@
Graph OOD Generalization: In this scenario, we assume that distribution shifts exist between source and target distributions, where $\mathcal{P}_{S}\neq\mathcal{P}_{T}$, and that $\mathcal{P}_{T}$ is not observable during model training. The objective of node-level graph OOD generalization is to learn an optimal model $\varphi_{{\theta}^{*}}$ using the nodes and labels from the source graph $(\mathcal{G}_{S},\textbf{Y}_{S})$, such that it generalizes well to the unseen target graph $\mathcal{G_{T}}$.
@@
If $\textbf{Y}_{T}=\varnothing$, indicating that none of the nodes in the target graph are labeled, the problem is classified as unsupervised domain adaptation. In this case, the learning objective $l(\cdot)$ relies solely on the labeled source domain data without direct supervision from the target domain. Otherwise, when a portion of the nodes in the target graph are labeled, the problem is considered semi-supervised, and $l(\cdot)$ can utilize both the source domain labels and the partial target domain labels to optimize model performance.
@@
Test-time Graph OOD Adaptation: In this scenario, a pre-trained model $\varphi_{\theta^{\prime}}$ and a target graph $\mathcal{G_{T}}$ with node labels $\textbf{Y}_{T}$ (which may be partially labeled or entirely unlabeled) are given. The goal is to adapt the pre-trained model during the test-time training phase in response to distribution shifts in the target domain. This process is performed without access to source domain data during adaptation. Given the distribution shifts between the source and target domains, the objective is to adjust the pre-trained model $\varphi_{\theta^{\prime}}$ at test time to optimize its performance on the target graph:
@@
Here, $\theta^{\prime}$ represents the parameters from the pre-trained model, while $\theta$ denotes the parameters introduced for test-time adaptation.
II-B Graph Distribution Shift Types
In traditional machine learning, several studies have discussed and defined various types of distribution shifts [25, 26], of which the most widely-used concepts are covariate shifts (i.e., shifts in $\mathcal{P}(\textbf{X})$) and concept shifts (i.e., shifts in $\mathcal{P}(\textbf{X}|\textbf{Y})$ or $\mathcal{P}(\textbf{Y}|\textbf{X})$). These concepts describe how data distributions change between training and test phases or across different domains. These shifts can naturally be extended to graph setting by replacing feature inputs X with the graph inputs $\mathcal{G}=(\textbf{A},\textbf{X})$.
@@ II-B Graph Distribution Shift Types
The impact of these shifts on graph learning tasks is multifaceted. Covariate shifts, particularly in graph structure, can lead to severe performance degradation, as models might not effectively learn the underlying relationships between nodes when their connectivity patterns differ. Concept shifts, on the other hand, often lead to misalignments in the model’s assumptions about the feature-label or structure-label dependencies. This issue is particularly problematic in tasks that heavily rely on the specific relationships between the graph components, such as link prediction or node classification.
II-C Discussion on Related Topics
Several topics are closely related to graph OOD learning, including: graph transfer learning, graph domain adaptation, graph OOD detection, and graph continual learning. While these topics share similar goals with graph OOD learning, they differ in scope and methodology. Below, we discuss their distinctions and connections.
Graph Transfer Learning. Traditionally, transfer learning is a more wide-ranging concept and involves transferring knowledge across distribution changes and diverse tasks. In contrast to the focus of this survey—graph learning under distribution shifts—graph transfer learning is not solely concerned with addressing distribution shifts. It also includes leveraging learned representations, model parameters, or structural patterns gained from one task to enhance performance in a different task.
Graph Domain Adaptation. Following traditional domain adaptation, graph domain adaptation methods [21, 22] typically rely on the covariate shift assumption, which posits an invariant relationship between graph inputs and labels, represented as $\mathcal{P}_{S}(\textbf{Y}|\textbf{A},\textbf{X})=\mathcal{P}_{T}(\textbf{Y}|\textbf{A},\textbf{X})$. The objective is to address distribution shifts in the input graph space across domains, i.e., $\mathcal{P}_{S}(\textbf{A},\textbf{X})\neq\mathcal{P}_{T}(\textbf{A},\textbf{X})$. In contrast to graph domain adaptation, graph OOD learning is more general and comprehensive, involving distribution shifts that go beyond the covariate shift assumption.
@@ II-C Discussion on Related Topics
II-D Proposed Taxonomy
As highlighted in the problem definitions, graph OOD generalization, training-time graph OOD adaptation, and test-time graph OOD adaptation, significantly differ in their learning settings. While these scenarios are related by their common focus on handling distribution shifts in graph learning, they represent different stages of dealing with distribution shifts: from generalization with no access to target data, to training-time adaptation where both domains are considered, to test-time adaptation when the pre-trained model and target domain data are available. Understanding these relationships helps form a more coherent view of how models can adapt to shifting data distributions. Accordingly, in the following three sections, we firstly categorize existing methods into graph OOD generalization, training-time graph OOD adaptation, and test-time graph OOD adaptation. Within each section, we further classify methods into model-centric and data-centric approaches. Model-centric approaches center on the learning process or the design of the graph model, while data-centric approaches emphasize the manipulation of input graphs, such as transforming graph structures or features.
Our taxonomy is illstrated in Figure 2. Note that even though the method categories for the three scenarios—graph OOD generalization, training-time graph OOD adaptation, and test-time graph OOD adaptation—may converge in terms of their high-level strategies (e.g., invariant representation learning, graph data augmentation), their application and emphasis on specific aspects, such as the data availability and the stage of the learning process, are what distinguish them.
III Graph OOD Generalization
Graph OOD generalization addresses the challenge of learning models that remain effective under distribution shifts, where the test data follows a different distribution from the training data [33, 34, 35, 36]. The setting of graph OOD generalization assumes that the model has no access to the target domain during training, making generalization particularly challenging [5, 18]. The goal is to develop models that can capture domain-invariant patterns and generalize across diverse, unseen target distributions without prior knowledge of the target domain’s characteristics [37]. In this section, we provide a comprehensive review of the most recent advances in graph OOD generalization, highlighting both model-centric and data-centric approaches. A summary of existing methods is presented in Table II.
III-A Model-Centric Approaches